### v2

# Rings and Boolean Algebras as Algebraic Theories

Abstract
We present a unified framework for representing commutative and Boolean rings through affine and hyperaffine algebraic theories. This yields categorical equivalences between these theories and rings, and leads to a new analysis of certain classes of modules over Boolean rings. The resulting structures naturally capture the algebraic semantics of the if-then-else construct in programming languages.
1 Introduction
An algebraic theory (also called a clone) is an abstraction of the notion of algebraic operations and the equations they satisfy. In classical universal algebra, one studies a variety by specifying a signature of operations and a set of equational identities. An algebraic theory encodes the same information in different terms: it is a collection of abstract operations of any finite arities together with a rule for composing them. An algebra of the theory is then a set equipped with actual operations corresponding to the elements of the theory, in such a way that composition is respected.
Given an algebraic structure, for example a ring $R$, a natural and intriguing question is how to represent $R$ as an algebraic theory. More precisely, we ask whether it exists a full embedding of the category of rings into that of algebraic theories. The first idea that comes to mind is to associate with $R$ the theory of (left) $R$-modules. The abstract operations, in this case, are the tuples $(r_{1},\ldots,r_{n})\in R^{n}$, thought of as linear combinations $r_{1}x_{1}+\cdots+r_{n}x_{n}$. If the ring is commutative, all the operations are commutative. Two binary operations $s,r$ are said to commute if
$r(s(x,y),s(z,w))=s(r(x,z),r(y,w))\text{,}$ (1)
and this condition can be generalised to higher arities. Despite this being a full embedding, if the goal is to reconstruct $R$ from abstract properties like (1), the encoding of $R$ via $R$-modules is not sufficiently rich.
It turns out that there are two orthogonal constructions that allow to reconstruct $R$ perfectly. The first, which works in the case of a commutative ring, associates with $R$ the theory of affine $R$-modules, namely the theory whose operations are linear combinations $r_{1}x_{1}+\cdots+r_{n}x_{n}$ such that $r_{1}+\cdots+r_{n}=1$. This theory further satisfies idempotence, which in the binary case takes the form
$r(x,x)=x\text{.}$ (2)
The intuition consists in imagining $r(x,y)$ as the affine combination rx+(1r)yrx+(1-r)y. The second construction, specific to Boolean rings, associates with a Boolean ring $B$ the theory of “hyperaffine” $B$-modules, that is, the theory whose operations are linear combinations $b_{1}x_{1}+\cdots+b_{n}x_{n}$ such that $b_{1}+\cdots+b_{n}=1$ and $b_{i}b_{j}=0$ for $i\neq j$. In this case, a strong property holds
$b(b(x,y),b(z,w))=b(x,w)\text{,}$ (3)
making each binary operation a rectangular band. What these two constructions have in common is the idea of encoding the ring through the binary operations of the corresponding theory. Given precise definitions of affine and hyperaffine theories, these observations can be crystallised in the following formal statement: for every commutative (Boolean) ring $R$, the theory of (hyper)affine $R$-modules is (hyper)affine; moreover, any (hyper)affine theory is uniquely determined by the ring of its binary operations (cf. Theorems 3.1 and 3.2).
One of the reasons that make this approach interesting is that studying the algebras of affine and hyperaffine theories provides an elegant aetiology for the axiomatisation of if-then-else construct in programming languages. For example, among the axioms for if-then-else originally introduced by McCarthy [27, p. 26], one corresponds to idempotence ((2) above), others encode the laws of composition and unitality in a clone, others express (3), and another one is commutativity (1). Given a Boolean ring $B$, a set together with an action of $B$ satisfying these conditions is what Bergman called a $B$-set [2]. It is natural to ask how these axioms could be weakened to capture alternative or more nuanced behaviors of the if-then-else operator.
Our first main contribution is the development of a convenient and uniform framework in which the “hyperaffine” and “affine” constructions can be realised. We provide a streamlined account of results in [16, 10], yielding a new proof that hyperaffine theories form a category equivalent to that of Boolean rings, as stated in [10, Theorem 4.9]. We then apply the same strategy to show that affine theories correspond precisely to commutative rings. Although some ingredients already appear in the literature, to the best of our knowledge there is no source where Theorem 3.2 is stated in the form adopted here; nor, more importantly, where both constructions are presented within a single, coherent framework. We believe that this unified perspective clarifies the parallels between the two settings and opens the way to further generalisations. Our second main contribution is the study of the models of an affine theory over a Boolean ring $B$. We provide two characterisations. The first, Theorem 4.1, is purely algebraic and parallels the definition of a $B$-set: the models of an affine theory over a Boolean ring are Boolean vector spaces equipped with a compatible action of $B$ which is idempotent and commutative. The second, Theorem 4.3, shows that these structures admit a sheaf representation analogous to $B$-sets (cf. Theorem 4.2). If $B$-sets provide the standard semantics for the if-then-else operator, we speculate that affine $B$-modules could be applied in a semantics of if-then-else in linear contexts, although a thorough investigation of this is left for future work.
1.0.1 Related works.
Hyperaffine theories were introduced by Johnstone [16] in the context of a theorem providing necessary and sufficient syntactic conditions for a variety of algebras to form a cartesian closed category. This result has since been revisited within an elegant and powerful framework, opening up new perspectives and developments [10, 11]. Among other contributions, [10] offers a detailed account of the connection between hyperaffine theories and their models, namely the so-called $B$-sets [2]. The study of these algebras had been further developed in [35, 36]. Moreover, hyperaffine theories are closely related to recent work in many-valued algebraic logic [6, 33, 5]. On the other hand, affine theories essentially capture what the Polish school calls Mal’cev modes [32], culminating a long line of research on varieties of affine modules [34, 7, 38, 31]. In turn, reducts of models of these theories are algebras known as convex spaces or barycentric algebras [37, 30, 20], which play a significant role in algebraic accounts of probability [18, 15]. The models of these theories, both affine and hyperaffine, are deeply connected with the algebraic treatment of the if-then-else construct, a topic that has been extensively studied in the semantics of programming languages [12, 29, 24, 25]. The clearest account of the semantic behaviour of if-then-else, as originally introduced by McCarthy [27], is arguably Bergman’s formulation [2], based on the notion of an action of a Boolean ring on a set. Alternative approaches have been proposed in which the Boolean ring $B$ acts not on sets but on other algebraic structures: join-semilattices with bottom [26], where programs are viewed as ordered structures, or semigroups [14], where programs are viewed as functions, thus creating a bridge with Kleene algebras with tests [21].
1.0.2 Outline of the paper.
Section 2 introduces the preliminary material on algebraic theories. Section 3 is divided into two main parts. In the first, it is shown that hyperaffine theories form a category equivalent to that of Boolean rings. In the second part, we apply the previously developed blueprint to show that affine theories, by contrast, correspond to commutative rings. Section 4 is devoted to the study of the models of these theories, with particular attention to the models of affine theories when the ring of coefficients is Boolean. Finally, Section 5 wraps up with concluding remarks and directions for future work.
2 Varieties and algebraic theories
In a broad sense, a variety is the category of models of a finitary and one-sorted equational theory. A given variety may be axiomatised by operations and equations in many ways; however, there is always a canonical choice, which is captured by the notion of algebraic theory due to Lawvere [22] (see also [3, Section 3] and [1, Chapter 11]). Here we give a formulation of algebraic theory essentially equivalent to what universal algebraists call an abstract clone [39].
Definition 1
An algebraic theory $T$ is a family of sets $T(n)$, for each $n\in\mathbb{N}$, together with elements $\pi^{n}_{i}\in T(n)$, $1\leq i\leq n$, and a composition operator $\circ:T(n)\times T(m)^{n}\to T(m)$, for each $m,n\in\mathbb{N}$. These data have to satisfy, writing $f(g_{1},\ldots,g_{n})$ in place of $f\circ(g_{1},\ldots,g_{n})$, two unit laws for $1\leq i\leq n$
$\pi^{n}_{i}(g_{1},\ldots,g_{n})=g_{i}\quad\text{ and }\quad f(\pi^{n}_{1},\ldots,\pi^{n}_{n})=f$ (4)
and the associative law
$f(g_{1}(h_{1},\ldots h_{m}),\ldots,g_{n}(h_{1},\ldots,h_{m}))=f(g_{1},\ldots,g_{n})\circ(h_{1},\ldots h_{m})\text{.}$ (5)
We call an element of $T(n)$ an $n$-ary operation in $T$.
Definition 2
A morphism of theories $\varphi:T\to T^{\prime}$ comprises, for every $n$, a function $\varphi:T(n)\to T^{\prime}(n)$ satisfying
$\varphi(\pi^{n}_{i})=\pi^{n}_{i}\text{ for }1\leq i\leq n\text{ and }\varphi(f(g_{1},\ldots,g_{n}))=\varphi(f)(\varphi(g_{1}),\ldots,\varphi(g_{n}))$ (6)
for all $f\in T(n)$ and $g_{1},\ldots,g_{n}\in T(m)$.
Definition 3
A model or an algebra for an algebraic theory $T$ is a set $X$ together with, for each $n\in\mathbb{N}$, a left action $\cdot:T(n)\times X^{n}\to X$ satisfying two conditions:
$\pi^{n}_{i}\cdot(x_{1},\ldots,x_{n})=x_{i}$ (7)
for all $1\leq i\leq n$, and
$f(g_{1},\ldots,g_{n})\cdot(x_{1},\ldots,x_{m})=f\cdot(g_{1}\cdot(x_{1},\ldots,x_{m}),\ldots,g_{n}\cdot(x_{1},\ldots,x_{m}))$ (8)
for each $f\in T(n),g_{1},\ldots,g_{n}\in T(m),x_{1},\ldots,x_{m}\in X$. The category of models of a theory is called a variety.
Given an algebraic theory $T$, the free model on $n$ generators, $n\in\mathbb{N}$, is given by $T(n)$ itself. There is an obvious forgetful functor from the category of models of $T$ to the category of sets, associating with the model $X$ its underlying set $X$. The underlying set of the free model on $n$ generators is $T(n)$. In this light, we can freely identify the operation $f\in T(n)$ with the element $f(x_{1},\ldots,x_{n})$ in the free model on the $n$ generators $x_{1},\ldots,x_{n}$.
The initial theory $S$ is such that $S(n)$ contains only the $n$ projections $\pi^{n}_{i}$ for $n\geq 1$ and $S(0)$ is empty; models of $S$ are mere sets. The terminal theory $U$ given by $U(n)=\{*\}$ for all $n$ and its only model is the singleton. The terminal theory has a subtheory $U^{\prime}$ given by
$U^{\prime}(n)=\begin{cases}\varnothing&\text{ if }n=0\\ \{*\}&\text{ if }n\neq 0\text{,}\end{cases}$
whose models are either the empty set or the singleton. We refer to $U$ and to $U^{\prime}$ as the two degenerate theories.
2.0.1 Mal’cev operations, affine and hyperaffine theories.
We recall the following well-known definitions.
Definition 4
Let $T$ be an algebraic theory; $f\in T(n)$ is said to be idempotent if $f(x_{1},\ldots,x_{1})=x_{1}$; $f\in T(n)$ and $g\in T(m)$ are said to commute if
$f(g(x^{1}_{1},\ldots x^{1}_{m}),\ldots,g(x^{n}_{1},\ldots x^{n}_{m}))=g(f(x^{1}_{1},\ldots x^{n}_{1}),\ldots,f(x^{1}_{m},\ldots x^{n}_{m}))\text{.}$
Moreover $f\in T(n)$ i said to split if
$f(f(x^{1}_{1},\ldots x^{1}_{n}),\ldots,f(x^{n}_{1},\ldots,x^{n}_{n}))=f(x^{1}_{1},\ldots,x^{n}_{n})\text{.}$
An idempotent, splitting operation is called a decomposition operation. A theory is idempotent (resp. commutative) if every operation is (resp. if every pair of operations commute).
A variety whose theory is idempotent and commutative is called a mode [32]. Another essential ingredient in the paper is that of a Mal’cev operation. Roughly speaking, a Mal’cev operation in a commutative theory generalises the operation f(x,y,z)=xy+zf(x,y,z)=x-y+z in Abelian groups. The importance of this concept lies in the fact that the presence of a Mal’cev operation in a theory is reflected in the internal structure of its models: given two congruences $R$, $S$ on a model $X$ of $T$, their composition $R\circ S$ commute, meaning that $R\circ S=S\circ R$, iff $T(3)$ has a Mal’cev operation (see [4, Theorem 2.2.2]). Formally, a Mal’cev operation is defined as follows; for an exhaustive and much more general treatment we refer the reader to the monograph [4].
Definition 5(see [4] Definition 2.2.1)
Let $T$ be an algebraic theory. An element $p\in T(3)$ is called a Mal’cev operation if
[…]
The following lemma is straightforward.
Lemma 1
Let $p\in T(3)$ be a Mal’cev operation that commutes with itself. Then $p$ satisfies
M1.
$p(x,y,p(t,u,v))=p(p(x,y,t),u,v)$;
M2.
$p(x,y,z)=p(z,y,x)$.
Moreover, (M1) holds iff
[…]
Finally, if $p$ and $q$ are two Mal’cev operations in $T(3)$ which commute with each other, then they coincide.
The category of algebraic theories is equivalent to the category of finitary monads on sets [23]; this equivalence commutes with the functors sending a monad to its algebras, and an algebraic theory to its models. Idempotent theories correspond to (finitary) monads on $\mathbf{Set}$ preserving the terminal object. Commutative theories have the desirable property that their categories of models admit the structure of symmetric monoidal closed category, and they correspond to (finitary) commutative monads on $\mathbf{Set}$ as defined by Kock [19].
Definition 6
We call an idempotent, commutative theory $T$ with a Mal’cev operation $p\in T(3)$ affine [^1] Note that this terminology is slightly non-standard and some authors use the word ‘affine’ to denote what we call idempotent theories.. Moreover, following [16, Section 4] we call an idempotent, commutative theory in which every operation splits hyperaffine.
The two degenerate theories $U$ and $U^{\prime}$ are affine and hyperaffine. In the equivalence between finitary monads and algebraic theories, hyperaffine theories correspond to those finitary monads on $\mathbf{Set}$ that preserve finite products, and neglecting the degenerate theory $U^{\prime}$, they correspond to monads preserving finite limits (see [17, Theorem 2.1] and subsequent discussion).
An easy calculation proves the following distributive property that will greatly simplify proofs concerning hyperaffine theories.
Lemma 2
Let $T$ be a hyperaffine theory. Then for $f\in T(k)$ and $g_{1},\ldots,g_{k}\in T(n)$, letting $t_{i}=f(x^{1}_{i},\ldots,x^{k}_{i})$ for $1\leq i\leq n$, we have
$f(g_{1}(t_{1},\ldots,t_{n}),\ldots,g_{k}(t_{1},\ldots,t_{n}))=f(g_{1}(x^{1}_{1},\ldots,x^{1}_{n}),\ldots,g_{k}(x^{k}_{1},\ldots,x^{k}_{n}))\text{.}$ (9)
Moreover, if an idempotent theory $T$ satisfies (9), then $T$ is hyperaffine.
3 Representig Boolean rings as algebraic theories
Any ring $R$ (assumed to be unital) gives rise to an algebraic theory $T_{R}$ as follows. For $n\in\mathbb{N}$, let $T_{R}(n)$ be the set of tuples $(r_{1},\ldots,r_{n})\in R^{n}$. These are the abstract $n$-ary operations of the theory. The projections in $T_{R}(n)$ are given by the standard basis of $R^{n}$: $(1,0,\ldots,0),\ldots,(0,\ldots,0,1)$, while composition $T_{R}(n)\times T_{R}(m)^{n}\to T_{R}(m)$ is given by matrix multiplication: $(A,B)\mapsto B\cdot A$. More about theories, called matricial, arising from matrices can be found in [9]. Models of the theory $T_{R}$ are precisely (left) modules over the ring $R$.
3.1 Hyperaffine theories
We start by examining the known case of hyperaffine theories.
Definition 7
For a given Boolean ring $B$, we define the algebraic theory $H_{B}$ whose operations in $H_{B}(n)$ are $(b_{1},\ldots,b_{n})\in B^{n}$ such that $b_{1}+\cdots+b_{n}=1$ and $b_{i}b_{j}=0$ for $i\neq j$.
As expected, a model of the theory $H_{B}$, is a $B$-module $X$ whose $n$-ary operations $f(x_{1},\ldots,x_{n})=b_{1}x_{1}+\cdots+b_{n}x_{n}$ with $b_{i}\in B$ and $x_{i}\in X$ satisfy $b_{1}+\cdots+b_{n}=1$ and $b_{i}b_{j}=0$ for $i\neq j$. Recall that Boolean rings can be equivalently described as Boolean algebras defining $a\land b:=ab$, $a\lor b=a+(1-a)b$, $\lnot a:=1-a$. Let $a,b\in B$; if $ab=a\land b=0$, then $a+b=a\lor b$. As a consequence, the operations in $H_{B}(n)$ are $(b_{1},\ldots,b_{n})\in B^{n}$ such that $b_{1}\lor\cdots\lor b_{n}=1$ and $b_{i}\land b_{j}=0$ for $i\neq j$.
Lemma 3
Let $B$ be a Boolean ring. The algebraic theory $H_{B}$ is hyperaffine.
Proof
As $B$ is commutative, $H_{B}$ is commutative. Idempotence is ensured by the condition $b_{1}+\cdots+b_{n}=1$. The fact that $b_{i}^{2}=b_{i}$ for all $i$ and $b_{i}b_{j}=0$ for $i\neq j$ precisely yields that every operation splits. ∎
Lemma 4(Proposition 4.4, [10])
Let $T$ be a hyperaffine theory. Then the set $T(2)$ is a Boolean algebra with operations:
  • $(a\land b)(x,y):=a(b(x,y),y)$;
  • $(a\lor b)(x,y):=a(x,b(x,y))$;
  • $(\lnot a)(x,y):=a(y,x)$;
  • $0(x,y):=y$ and $1(x,y):=x$.
Proof
The only nontrivial identity to prove is $(a\land b)\lor(a\land c)=a\land(b\lor c)$. It can be obtained, through some calculations, as a consequence of Lemma 2 in the form: $a(b(a(x,y),a(z,w)),c(a(x,y),a(z,w)))=a(b(x,z),c(y,w))\text{.}$
When the Boolean ring $B$ is degenerate, that is to say when $0=1$, we obtain the degenerate theory $U$. There is an almost perfect correspondence between hyperaffine theories and Boolean rings. This correspondence is imperfect in the sense that the other degenerate theory $U^{\prime}$ does not come from a ring.
We recall the following notion introduced by Johnstone [16, p.461].
Definition 8
Let $T$ be an algebraic theory, and let $f\in T(n)$. We call the following elements of $T(2)$ coefficients of $f$:
$f[1](x,y):=f(x,y,\ldots,y),\ldots,f[n](x,y):=f(y,\ldots,y,x)\text{.}$
The next proposition shows in particular the fact that any operation in a hyperaffine theory is completely determined by its coefficients.
Proposition 1
Let $T$ be a hyperaffine theory and let $B$ be the Boolean ring $T(2)$. The following hold:
  1. 1.
    if $f\in T(n)$, then $f[1]\lor\cdots\lor f[n]=1$ and $f[i]f[j]=0$ for $i\neq j$;
  2. 2.
    for $f,g\in T(n)$, if $f[i]=g[i]$ for all $1\leq i\leq n$, then $f=g$;
  3. 3.
    for every $(b_{1},\ldots,b_{n})$ such that $b_{1}\lor\cdots\lor b_{n}=1$ and $b_{i}b_{j}=0$ for $i\neq j$, there is $f\in T(n)$ such that $f[i]=b_{i}$ for all $1\leq i\leq n$.
Proof
The first item is easy. The second item is [10, Lemma 4.6(i)]. The third item can be deduced from the proof of [10, Proposition 4.7]. ∎
We present here a new proof of (part of) [10, Theorem 4.9].
Theorem 3.1
The full subcategory of hyperaffine theories different than $U^{\prime}$ is equivalent to the category of Boolean rings.
Proof
Let $H\neq U^{\prime}$ be a hyperaffine theory, and let $B$ be a Boolean ring. The two assignments $H\mapsto H(2)$ and $B\mapsto H_{B}$ are functorial. It is easy to see that $B$ is isomorphic to $H_{B}(2)$. For each $n\in\mathbb{N}$, we consider $\varphi:H(n)\to H(2)^{n}$ given by $\varphi(f)=(f[1],\ldots,f[n])$. By Proposition 1, $\varphi$ is a bijection onto the image $\{(b_{1},\ldots,b_{n})\in H(2)^{n}:b_{1}\lor\cdots\lor b_{n}=1,r_{i}r_{j}=0\}$. We prove that for every $f\in H(n)$ and $g_{1},\ldots,g_{n}\in H(m)$
$\varphi(f)(\varphi(g_{1}),\ldots,\varphi(g_{n}))(x_{1},\ldots,x_{m})=\varphi(f(g_{1},\ldots,g_{n}))(x_{1},\ldots,x_{m})\text{.}$
By definition of $\varphi$ this amounts to prove that
$\begin{bmatrix}g_{1}[1]&\cdots&g_{n}[1]\\ \vdots&\ddots&\vdots\\ g_{1}[m]&\cdots&g_{n}[m]\end{bmatrix}\begin{bmatrix}f[1]\\ \vdots\\ f[n]\end{bmatrix}(x,y)=\begin{bmatrix}f(g_{1},\ldots,g_{n})[1]\\ \vdots\\ f(g_{1},\ldots,g_{n})[m]\end{bmatrix}(x,y)\text{,}$
i.e. that for every $i=1,\ldots,m$,
$f[1]g_{1}[i](x,y)\lor\cdots\lor f[n]g_{n}[i](x,y)=f(g_{1}[i](x,y),\ldots,g_{n}[i](x,y))\text{.}$
Through some calculations, employing idempotence of $f$ and Lemma 2, one can obtain the result inductively. ∎
3.1.1 Hyperaffine theories and algebraic logic.
In [8] it is shown that the variety of Boolean algebras can be presented using two nullary operations, 0 and $1$, and a ternary operation $\alpha$ called ‘conditional disjunction’. Expressed in terms of the Boolean operators, conditional disjunction is: $\alpha(a,x,y)=(a\land x)\lor(\lnot a\land y)$, i.e. if $a$ then $x$ else $y$. Restricting the conditions defining hyperaffine theories (cf. Definition 6) to binary operations, Lemma 3 and Lemma 4 imply that the following is an axiomatisation of Boolean algebras in the signature with a ternary operation $\alpha$ (whose intended meaning is if-then-else) and two constants $0,1$:
H1.
$\alpha(a,1,0)=a$ cf. (4)
H2.
$\alpha(1,a,b)=a$ and $\alpha(0,a,b)=b$ cf. (4)
H3.
$\alpha(a,\alpha(b,x,y),\alpha(c,x,y))=\alpha(\alpha(a,b,c),x,y)$ cf. (5)
H4.
$\alpha(a,x,x)=x$ cf. Definition 6
H5.
$\alpha(a,\alpha(b,x,y),\alpha(b,z,w))=\alpha(b,\alpha(a,x,z),\alpha(a,y,w))$ cf. Definition 6
H6.
$\alpha(a,\alpha(a,x,y),\alpha(a,z,w))=\alpha(a,z,w)$ cf. Definition 6
Recently, these axioms have been adapted [33] from the two-dimensional case (featuring two constants, 0 and $1$, and a ternary if-then-else) to the $n$-dimensional case, with $n$ constants $0,\ldots,n-1$ acting as projections and a generalised $(n+1)$-ary if-then-else. The axiomatisation presented in [33, Definition 14] is equivalent to the $(n+1)$-ary form of axioms (H1)-(H6), thanks to Lemma 2. These algebras, called $n$-Boolean algebras (nnBAs) by the authors, have inspired subsequent work in algebraic logic aimed at generalising classical propositional logic to the case of an arbitrary finite number of symmetric truth values [33, 5]. The connection with hyperaffine theories is the following:
Proposition 2
If $T$ is a hyperaffine theory, then just as $T(2)$ is a Boolean algebra, $T(n)$ is an nnBA. Conversely, any sequence $(A_{n})_{n}$ of nnBAs sharing the same Boolean algebra of coordinates [6, Definition 10] determines a hyperaffine theory.
3.2 Commutative rings and affine theories
In the case of affine theories, the assumption that every operation is a decomposition operation (cf. Definition 4) is exchanged for the requirement of the existence of a Mal’cev operation. Remark that the only model of a Mal’cev operation $p$ that splits is degenerate:
$y=p(y,x,x)=p(p(x,x,y),p(x,y,y),p(x,y,y))=p(x,y,y)=x\text{,}$
and therefore the only hyperaffine theory that are also affine are the degenerate ones.
Definition 9
For a given commutative ring $R$, we define the algebraic theory $A_{R}$ whose operations in $A_{R}(n)$ are $(r_{1},\ldots,r_{n})\in R^{n}$ such that $r_{1}+\cdots+r_{n}=1$.
Lemma 5
The algebraic theory $A_{R}$ is idempotent and commutative. Moreover, the operation $(1,-1,1)\in A_{R}(3)$ is a Mal’cev operation.
Proof
It can be easily checked that the commutativity of the monoid $(R,\cdot,1)$ ensures that $A_{R}$ is commutative, and the condition $r_{1}+\cdots+r_{n}=1$ that $A_{R}$ is idempotent. ∎
The following lemma is a known result in universal algebra [32, Theorem 6.3.3]. The proof crucially depends on Lemma 1.
Lemma 6
Let $T$ be an affine theory. Then $T(2)$ endowed with the following structure:
  • $(a+b)(x,y):=p(a(x,y),y,b(x,y))$;
  • $(a\cdot b)(x,y):=a(b(x,y),y)$;
  • a(x,y):=p(y,a(x,y),y)-a(x,y):=p(y,a(x,y),y);
  • $0(x,y):=y$ and $1(x,y):=x$
is a commutative ring.
From this we infer the following important property.
Lemma 7
If $T$ be an affine theory, then
$p(a(x,y),b(x,y),c(x,y))=(a-b+c)(x,y)\text{.}$
Starting from a ring, applying both constructions returns an isomorphic ring.
Lemma 8
Let $R$ be a commutative ring. The ring $A_{R}(2)$ of Lemma 6 is isomorphic to $R$.
Proof
Recall that $A_{R}(n)=\{(r_{1},\ldots,r_{n})\in R^{n}:r_{1}+\cdots+r_{n}=1\}$. In particular $A_{R}(2)=\{(r,1-r):r\in R\}$. Then, by definition:
$(r,1-r)+(s,1-s)$ $=\begin{bmatrix}r&0&s\\ 1-r&1&1-s\end{bmatrix}\begin{bmatrix}1\\ -1\\ 1\end{bmatrix}=(r+s,1-r-s)$
$(r,1-r)\cdot(s,1-s)$ $=\begin{bmatrix}s&0\\ 1-s&1\end{bmatrix}\begin{bmatrix}r\\ 1-r\end{bmatrix}=(rs,1-rs)$
$-(r,1-r)$ $=\begin{bmatrix}0&r&0\\ 1&1-r&1\end{bmatrix}\begin{bmatrix}1\\ -1\\ 1\end{bmatrix}=(-r,1+r)$
This concludes the proof. ∎
Now, we prove the analogue of Proposition 1.
Proposition 3
Let $T$ be an affine theory. Let $R$ be the commutative ring $T(2)$. The following hold:
  1. 1.
    if $f\in T(n)$, then $f[1]+\cdots+f[n]=1$ in $R$;
  2. 2.
    for $f,g\in T(n)$, if $f[i]=g[i]$ for all $1\leq i\leq n$, then $f=g$;
  3. 3.
    for every $(r_{1},\ldots,r_{n})\in R^{n}$ such that $r_{1}+\cdots+r_{n}=1$, there is $f\in T(n)$ such that $f[i]=r_{i}$ for all $1\leq i\leq n$.
Proof
Regarding the first item, we prove that $f[1]+\cdots+f[n]=1$. Observe that
$(f[1]+f[2])(x,y)$ $=p(f[1](x,y),y,f[2](x,y))$
$=p(f(x,y,\ldots,y),f(y,y\ldots,y),f(y,x,\ldots,y))$ $f\text{ is idemp.}$
$=f(p(x,y,y),p(y,y,x),\ldots,p(y,y,y))$ $p,f\text{ commute}$
$=f(x,x,y\ldots,y)$ $p\text{ is Mal'cev}$
and, therefore, inductively,
$(f[1]+\cdots+f[n])(x,y)=f(x,x,\ldots,x)=x\text{.}$
For the second item, observe that
$p(f[1](x_{1},y),y,f[2](x_{2},y))$ $=p(f(x_{1},y,\ldots,y),f(y,y\ldots,y),f(y,x_{2},\ldots,y))$
$=f(p(x_{1},y,y),p(y,y,x_{2}),\ldots,p(y,y,y))$
$=f(x_{1},x_{2},y\ldots,y)$
and, therefore, inductively,
$p(f(x_{1},\ldots,x_{n-1},y),y,f[n](x_{n},y))=f(x_{1},x_{2},\ldots,x_{n})\text{.}$
This implies that if $f[i]=g[i]$ for all $i=1,\ldots,n$, then $f=g$.
Finally, we prove the third item. The proof is by induction on $n$. If $n=1$, then clearly $f(x)=x$. Assume that the result holds for n1n-1, and let $(r_{1},\ldots,r_{n})\in R^{n}$ be such that $r_{1}+\cdots+r_{n}=1$. By inductive assumption there is $g\in T(n-1)$ such that $g[1]=r_{1}+r_{2}$ and $g[i]=r_{i+1}$ for $2\leq i\leq n$. Let $f(x_{1},\ldots,x_{n}):=p(r_{2}(x_{2},x_{1}),x_{1},g(x_{1},x_{3},\ldots,x_{n}))$. We prove that $f[i]=r_{i}$; we separate three cases $i=1,2$ and $3\leq i\leq n$:
$f(x,y,\ldots,y)$ $=p(r_{2}(y,x),x,g(x,y,\ldots,y))$
$=r_{2}(y,x)-x+g[1](x,y)$ by Lemma 7
$=(1-r_{2}-1+r_{1}+r_{2})(x,y)$ by inductive assumption
$=r_{1}(x,y)$
$f(y,x,y,\ldots,y)$ $=p(r_{2}(x,y),y,g(y,\ldots,y))$
$=p(r_{2}(x,y),y,y)$ $g\text{ is idempotent}$
$=r_{2}(x,y)$ $p\text{ is Mal'cev}$
$f[i](x,y)$ $=p(r_{2}(y,y),y,g[i-1](x,y))$ $3\leq i\leq n$
$=p(y,y,g[i-1](x,y))$ $r_{2}\text{ is idempotent}$
$=g[i-1](x,y)$ $p\text{ is Mal'cev}$
$=r_{i}(x,y)$ by inductive assumption
This concludes the proof. ∎
We have all the ingredients to present one of the main theorems.
Theorem 3.2
The full subcategory of affine theories different than $U^{\prime}$ is equivalent to the category of commutative rings.
Proof
Let $A\neq U^{\prime}$ be an affine theory, and let $R$ be a commutative ring. The two assignments $A\mapsto A(2)$ and $R\mapsto A_{R}$ are functorial. Lemma 8 says that $R$ is isomorphic to $A_{R}(2)$. For each $n\in\mathbb{N}$, consider the function $\varphi:A(n)\to A(2)^{n}$ such that $\varphi(f)=(f[1],\ldots,f[n])$. By Proposition 3, $\varphi$ is a bijection onto the image $\{(r_{1},\ldots,r_{n})\in A(2)^{n}:r_{1}+\cdots+r_{n}=1\}$. We prove that $\varphi$ is a morphism of theories, i.e. that for every $f\in A(n)$ and $g_{1},\ldots,g_{n}\in A(m)$
$\varphi(f)(\varphi(g_{1}),\ldots,\varphi(g_{n}))(x_{1},\ldots,x_{m})=\varphi(f(g_{1},\ldots,g_{n}))(x_{1},\ldots,x_{m})\text{.}$
By definition of $\varphi$ this amounts to prove that
$\begin{bmatrix}g_{1}[1]&\cdots&g_{n}[1]\\ \vdots&\ddots&\vdots\\ g_{1}[m]&\cdots&g_{n}[m]\end{bmatrix}\begin{bmatrix}f[1]\\ \vdots\\ f[n]\end{bmatrix}(x,y)=\begin{bmatrix}f(g_{1},\ldots,g_{n})[1]\\ \vdots\\ f(g_{1},\ldots,g_{n})[m]\end{bmatrix}(x,y)\text{,}$
i.e. that for every $i=1,\ldots,m$,
$f[1]g_{1}[i](x,y)+\cdots+f[n]g_{n}[i](x,y)=f(g_{1}[i](x,y),\ldots,g_{n}[i](x,y))\text{.}$
This is easily proved by induction. ∎
4 Boolean actions and the if-then-else construct
In this section we take a closer look at the models of hyperaffine theories, and affine theories whose ring of coefficients is Boolean. We begin with the following observation.
Lemma 9
Let $B$ be a Boolean ring. If $X$ is a model of $A_{B}$ or a model of $H_{B}$, then $X$ is endowed with a binary action of $B$, i.e. a function $B\times X^{2}\to X$, satisfying the following axioms:
R1.
$b(x,x)=x$;
R2.
$b(c(x,y),c(z,w))=c(b(x,z),b(y,w))$;
R3.
$1(x,y)=x$ and $0(x,y)=y$;
R4.
a(b(x,y),c(x,y))=(ab+(1a)c)(x,y)a(b(x,y),c(x,y))=(ab+(1-a)c)(x,y)
for all $x,y,z,w\in X$ and $a,b,c\in R$.
Proof
Note that (R1) follows from the idempotence of $A_{B}$ (or $H_{B}$), (R2) from the commutativity, while (R3) and (R4) from the fact that $X$ is a model (cf. conditions (7) and (8) of the first section). ∎
Remarkably, a set of axioms similar to (R1)-(R4) was given by Bergman with the purpose of capturing the concept of action of $B$ on a set.
Definition 10(Definition 7, [2])
Let $B$ be a Boolean ring. A $B$-set is a set $X$ together with a function $b:X^{2}\to X$ for each $b\in B$ satisfying the following requirements:
B1.
$b(x,x)=x$;
B2.
$b(b(x,y),z)=b(x,z)=b(x,b(y,z))$;
B3.
(1b)(x,y)=b(y,x)(1-b)(x,y)=b(y,x);
B4.
$0(x,y)=y$;
B5.
$b(c(x,y),y)=bc(x,y)$
for all $x,y\in X$ and $b,c\in B$.
Remark 1
More generally, an action of $B$ on an algebra is an action on the underlying set of the algebra, such that each operation $b\in B$ commutes with each operation of the algebra.
Firstly, we establish that (B1)-(B5) can be rewritten as (R1)-(R4) plus an additional axiom.
Lemma 10
For a Boolean ring $B$, (B1)-(B5) are equivalent to (R1)-(R4) plus
R5.
$b(b(x,y),b(z,w))=b(x,w)$
for all $x,y,z,w\in X$ and $b\in B$.
Proof
Assume (R1)-(R5). (R5) combined with (R1) immediately gives (B2). (R4) and (R3) give (B3). (B5) follows from (R4) with $q=0$. Conversely, assume (B1)-(B5). The content of [2, Proposition 11] is precisely that (B1)-(B5) imply (R2). (R3) is obvious in light of (B3) and (B4). It is not difficult to see that (R5) follows repeatedly applying (B2). To obtain (R4), first observe that $ab+(1-a)c=ab\lor(1-a)c$ as ab(1a)c=0ab(1-a)c=0. Moreover, it easy to see that
$(a\lor b)(x,y)=a(x,b(x,y))\text{.}$ (10)
Therefore:
[…]
$=a(a(b(x,y),b(x,c(x,y))),a(y,c(x,y)))$ (R2)
$=a(b(x,y),c(x,y))$ (R5)
This concludes the proof. ∎
Algebras for a hyperaffine theory $H$ are exactly sets equipped with the Boolean action of $H(2)$. We present a conceptually simpler (as it does not rely on Bergman’s Theorem 4.2) proof of [10, Proposition 3.16].
Proposition 4
Let $H$ be a hyperaffine theory, and let $B:=H(2)$ be the Boolean ring associated with $H$. Then the models of $H$ are sets equipped with an action of $B$ satisfying (R1)-(R5), or, equivalently by Lemma 10, $B$-sets according to Definition 10.
Proof
The category of $B$-sets is clearly a variety, with algebraic theory $T$ generated by the binary operations $\{b(x,y):b\in B\}$ modulo the axioms (R1)-(R5). The theory $T$ is idempotent by (R1), commutative by (R2), and every operation splits by (R5). By Theorem 3.1 it is therefore of the form $H_{B^{\prime}}$ for some Boolean ring $B^{\prime}$, and we are done if $B\simeq B^{\prime}$. We know that $B^{\prime}\simeq T(2)$. Let $f(x,y)$ be an element of $T(2)$. By (R3) and (B3) we can assume that $f(x,y)$ is the composition of binary operations of the form $d(x,y)$, with $d\in B$. We then apply (R4) in the form of the rewriting rule
$a(b(x,y),c(x,y))\to(ab+(1-a)c)(x,y)$
to get a unique normal form $b(x,y)$ with $b\in B$. This shows that $T(2)\simeq B$. ∎
Unlike the case of hyperaffine theories, binary operations do not completely determine an affine theory, due to the presence of the ternary Mal’cev operation $p$. However, axioms (R1)-(R4) can describe models of an affine theory if the Mal’cev operation can be expressed as composition of binary operations, and this holds iff there are no ring homomorphisms from $R$ to $\mathbb{F}_{2}$, where $\mathbb{F}_{2}$ is the field with two elements (see e.g. [38, 13]). In this case, models of an affine theory $A$ are sets with a binary action of the ring $R:=A(2)$ satisfying (R1)-(R4).
Axioms in the general case can be found in [32, Theorem 6.3.4]: for an affine theory $A$, and associated ring $R:=A(2)$, models of $A$ are sets equipped with binary operations $a:X^{2}\to X$ for each $a\in R$ satisfying (R1)-(R4) and a ternary operation $p:X^{3}\to X$ satisfying
A0.
$p$ is a Mal’cev operation;
A1.
$p$ is idempotent;
A2.
$p$ commutes with itself and with every $a\in R$;
A3.
$p(y,x,y)=(1+1)(y,x)$;
A4.
p(a(x,y),b(x,y),c(x,y))=(ab+c)(x,y)p(a(x,y),b(x,y),c(x,y))=(a-b+c)(x,y).
When $B$ is a Boolean ring, we obtain the following new characterisation.
Theorem 4.1
Let $B$ be a Boolean ring, and consider the affine theory $A_{B}$. Models of $A_{B}$, together with the choice of $o\in X$, are precisely vector spaces over $\mathbb{F}_{2}$ equipped with an action of $B$ satisfying (R1)-(R4) and such that
L1.
$b(x,y)+b(z,w)=b(x+z,y+w)$;
for every $a,b\in B$ and $x,y,z,w\in X$.
Proof
If $X$ models $A_{B}$, then $X$ is endowed with an action of $B$ and has a ternary operation satisfying (R1)-(R4) and (A0)-(A4). We endow $X$ with the structure of $\mathbb{F}_{2}$-vector space as follows. We define $x+y:=p(x,o,y)$; this sum is associative and commutative by (A1), (A2) and Lemma 1. The element $o\in X$ is a zero for the sum by (A0). Moreover, the inverse of $x$ is given by $x$ itself as (A3) shows:
$x+x=p(x,o,x)=0(x,o)=o\text{.}$
Then, we add the action of $\mathbb{F}_{2}$: $1x:=x$, $0x:=o$. The only nontrivial axiom is proven using (A3):
$(1+1)x=0x=o=0(x,o)=(1+1)(x,o)=p(x,o,x)=x+x=1x+1x\text{.}$
Finally, using (A2) it is easy to derive (L1). Conversely, given an $\mathbb{F}_{2}$-vector space $(X,+,0)$ with an action of $B$ that satisfies (R1)-(R4) and (L1), we define $p(x,y,z)=x+y+z$. As $X$ has characteristic $2$, (A0) and (A1) are verified. (A3) follows by definition of $p$:
$p(y,x,y)=y+x+y=(1+1)y+x=x=0(y,x)=(1+1)(y,x)\text{.}$
Moreover, (A2) follows from (L1). Now, observe that
$(a+b)(x,y)$ $=a(b(y,x),b(x,y))$ (R4),(B3)
$=a(b(x,y)+b(y,x)+b(x,y),0+b(x,y))$
$=a(b(x,y)+b(y,x),0)+b(x,y)$ (L1),(R1)
$=a(b(x+y,y+x),0)+b(x,y)$ (L1)
$=a(b(x+y,x+y),0)+b(x,y)$
$=a(x+y,0)+b(x,y)$ (L1),(R1)
$=a(x+y,y+y)+b(x,y)$
$=a(x,y)+y+b(x,y)$ (L1),(R1)
and therefore
$(a+b+c)(x,y)$ $=a(b(c(x,y),c(y,x)),b(c(y,x),c(x,y)))$ (R4)
$=a(b(y,x)+x+c(y,x),b(x,y)+y+c(x,y))$
$=a(b(y,x),b(x,y))+a(x,y)+a(c(y,x),c(x,y))$ (L1)
$=a(x,y)+y+b(x,y)+a(x,y)+a(x,y)+y+c(x,y)$
$=a(x,y)+b(x,y)+c(x,y)$
proving (A4) and concluding the proof. ∎
In [26], the semantics of if-then-else is given in terms of an action of $B$ on a $\lor$-semilattice with $\bot$, in the sense of Remark 1. Theorem 4.1 suggests that it might be interesting to investigate a semantics of if-then-else as a weakened (omitting (R5)) linear action of $B$ on an $\mathbb{F}_{2}$-vector space.
4.1 Boolean actions and sheaves
We now set out to give a sheaf representation of these structures. We first recall Bergman’s result. For a set $X$, we denote by $\operatorname{Equiv}(X)$ the set of equivalence relations on $X$ and let $\Delta_{X}:=\{(x,x):x\in X\}$ and $\nabla_{X}:=X^{2}$. An element $R\in\operatorname{Equiv}(X)$ is called a factor relation if there is $\lnot R\in\operatorname{Equiv}(X)$ such that $R\cap\lnot R=\Delta_{X}$ and $R\circ\lnot R=\nabla_{X}\text{.}$ It is well-known (cf. [28, Section 4.4]) that for a set $X$, there is a bijection between (i) binary operations $a:X^{2}\to X$ satisfying (R1) and (R5); (ii) factor relations on $X$; (iii) the decompositions of $X$ into two factors: $X\simeq A\times B$. Bergman’s next result can be seen as an enrichment of the foregoing fact. We recall that a sheaf $S:B^{\operatorname{op}}\to\mathbf{Set}$ is a functor $S$ (that is, a set $S(b)$ for every $b\in B$, and a function $S(b)\to S(c)$ when $c\leq b$) such that $S(b)\simeq S(c)\times S(d)$ when $b=c\lor d$ and $cd=0$.
Theorem 4.2(Lemma 10, Theorem 12, Lemma 16 [2])
Let $X$ be a set. A set $F$ of binary operations on $X$ satisfying (R1), (R2) and (R5) for all $b,c\in F$ gives rise to a Boolean ring $B$ and a $B$-set structure on $X$. Moreover, a $B$-set structure on $X$ is equivalently given by
  1. 1.
    a family of equivalence relations $\{R_{b}:b\in B\}\subseteq\operatorname{Equiv}(X)$ satisfying:
    1. (a)
      $R_{b}\circ R_{1-b}=\nabla_{X}$;
    2. (b)
      $b\leq c\implies R_{c}\subseteq R_{b}$;
    3. (c)
      $R_{b}\cap R_{c}\subseteq R_{b\lor c}$;
    4. (d)
      $R_{1}=\Delta_{X}$
  2. 2.
    a sheaf $S:B^{\operatorname{op}}\to\mathbf{Set}$ of sets on $B$ such that $X\simeq S(1)$.
Remark 2
More generally, for any variety of algebras $\mathcal{V}$, there is an equivalence between the category of algebras with action of a Boolean ring $B$, and the category of sheaves $B^{\operatorname{op}}\to\mathcal{V}$.
Thanks to Proposition 4.1 we obtain the following new representation for models of an affine theory whose ring of coefficients is Boolean.
Theorem 4.3
Let $B$ be a Boolean ring, and let $A_{B}$ its corresponding affine theory. Pointed models of $A_{B}$, that is models $X$ of $A_{B}$ together with the choice of $o\in X$, correspond precisely to sheaves $S:B^{\operatorname{op}}\to\mathbf{Vect}_{\mathbb{F}_{2}}$ of $\mathbb{F}_{2}$-vector spaces over $B$ such that $X\simeq S(1)$.
Proof
Firstly, observe that, by Remark 2, a sheaf $S:B^{\operatorname{op}}\to\mathbf{Vect}_{\mathbb{F}_{2}}$ can be equivalently described as a $\mathbb{F}_{2}$-vector space $X$, together with an action of $B$ satisfying (B1)-(B5) and (L1). Given a pointed model $(X,o)$ of $A_{B}$, we define an $\mathbb{F}_{2}$-vector as in Theorem 4.1. Now, $X$ has an action of $B$ that satisfies (R1)-(R4) and (L1) by Theorem 4.1. Using the usual algebraic manipulations involving the Mal’cev operation, we can see that:
$b(x,y)=b(x,o)+(1-b)(y,o)\text{.}$ (11)
This implies that the action satisfies (R5) as well (here we prove part of (B2)):
$b(b(x,y),z)$ $=b(b(x,o)+(1-b)(y,o))+(1-b)(z,o)$ (11)
$=b(b(x,o),o)+b((1-b)(y,o),o)+(1-b)(z,o)$ (L1)
$=b(x,o)+0(x,o)+(1-b)(z,o)$ (B5)
$=b(x,z)$ (11)
The converse is obvious in light of Theorem 4.1. ∎
5 Conclusions and vistas
This work has shown that affine and hyperaffine algebraic theories form a natural setting for representing commutative and Boolean rings, respectively. By developing a uniform framework that accommodates both constructions, we have clarified the features that make affine theories the counterpart of commutative rings, and hyperaffine theories the counterpart of Boolean rings. Moreover, the study of the models of both reveals a close connection with the algebraic treatment of the if-then-else construct. Several directions arise. A first step is to generalise the correspondence established here beyond rings, toward suitable classes of commutative semirings. A second line of investigation concerns the algebras described in Theorem 4.1, which suggest an equational foundation for an if-then-else operator in linear contexts, where the combination of programs obeys algebraic rather than set-theoretic structure. Finally, a deeper understanding of the free algebras of affine theories, might lead, inspired by [6], to an extension of the classical theory of affine combinations into higher dimensions.